### v2

# Summarization Metrics for Spanish and Basque: Do Automatic Scores and LLM-Judges Correlate with Humans?

Abstract
Studies on evaluation metrics and LLM-as-a-Judge models for automatic text summarization have largely been focused on English, limiting our understanding of their effectiveness in other languages. Through our new dataset BASSE (BAsque and Spanish Summarization Evaluation), we address this situation by collecting human judgments on 2,040 abstractive summaries in Basque and Spanish, generated either manually or by five LLMs with four different prompts. For each summary, annotators evaluated five criteria on a 5-point Likert scale: coherence, consistency, fluency, relevance, and 5W1H. We use these data to reevaluate traditional automatic metrics used for evaluating summaries, as well as several LLM-as-a-Judge models that show strong performance on this task in English. Our results show that currently proprietary judge LLMs have the highest correlation with human judgments, followed by criteria-specific automatic metrics, while open-sourced judge LLMs perform poorly. We release BASSE [^1] BASSE is available under open licenses at https://github.com/hitz-zentroa/summarization. and our code publicly, along with the first large-scale Basque summarization dataset containing 22,525 news articles with their subheads.
1 Introduction
Automatic text summarization seeks to create a concise and fluent text that captures the main ideas from one or more documents Giarelis, Mastrokostas, and Karacapilidis (2023). This task has been studied in depth in Natural Language Processing (NLP) Dang (2006); Dang and Owczarzak (2008) and plays an important role in mitigating excess information in an age of increasing information glut Navas-Martin, Albornos-Muñoz, and Escandell-García (2012).
Although extractive approaches, where sentences and phrases are taken directly from the original document, were originally more common McDonald (2007); Lin and Bilmes (2011), current approaches instead generally rely on abstractive methods, where the model generates novel sentences to capture the most salient information. This change in modeling paradigm also leads to new challenges for evaluation, as abstractive models are more prone to hallucinating information Maynez et al. (2020); Vasilyev, Dharnidharka, and Bohannon (2020); Laban et al. (2022).
In order to evaluate automatic summaries, researchers generally rely on three main strategies: i) human evaluation, ii) evaluation using reference summaries and automatic metrics, and iii) evaluation via Large Language Models (LLMs), often referred to as LLM-as-a-Judge. Human evaluation over a small number of important criteria, e.g., coherence or fluency, is often chosen to give high-quality evaluation of abstractive summaries. However, this approach is slow, expensive, and does not lend itself to internal evaluations while iterating over models. Therefore, most research instead relies on automatic metrics, e.g., ROUGE Lin (2004) and, more recently, LLM-as-a-Judge Leiter et al. (2023); Kim et al. (2024); Lyu et al. (2025).
Selecting appropriate metrics or judge LLMs relies on meta-evaluations Bhandari et al. (2020); Fabbri et al. (2021); Deutsch, Dror, and Roth (2022) that measure the correlation between human evaluations and automatic evaluation methods. While these studies provide invaluable information, they have only been performed comprehensively for English, leaving the assessment of their performance in other languages to be explored.
We therefore address this gap by collecting human annotations for automatic summaries in Basque and Spanish—languages with respective fragmentary and medium coverage of language resources Giagkou et al. (2023). Specifically, we collect abstractive summaries generated by five LLMs and four different prompts on 45 news articles in each language. We then collect human judgments for each summary with respect to five criteria—namely, coherence, consistency, fluency, relevance, and 5W1H. With this novel dataset, BAsque and Spanish Summarization Evaluation (BASSE), we then proceed to address the following research questions:
  • RQ1: Do automatic metrics for evaluating summarization correlate well with human judgments in Basque and Spanish?
  • RQ2: Do judge LLMs have the ability to replicate human judgments for summarization?
  • RQ3: Which summary criteria (coherence, consistency, fluency, relevance, and 5W1H) do state-of-the-art LLMs struggle with in Basque and Spanish?
Our results indicate that proprietary LLM-as-a-Judge models currently have the highest correlation with human judgments on our data, while several traditional metrics also correlate well with specific criteria. On the contrary, open-source judge LLMs currently do not correlate well with human judgments on our data.
Our contributions are hence the following:
  • We present a new multilingual corpus (Basque, Spanish) of summaries generated by five LLMs manually annotated for five criteria on a 5-point Likert-scale;
  • We study the correlation between human judgments with the results of automatic evaluation metrics and LLM judges in languages other than English;
  • We additionally provide the first corpus for summarization in Basque, consisting of 22,525 news articles and their subheads.
2 Related Work
Two main approaches exist in summarization: extractive and abstractive El-Kassas et al. (2021). The extractive approach selects and takes the most relevant sentences directly from the input document, concatenating them to create the summary. On the other hand, the abstractive approach generates a set of novel sentences to represent the main information of the original document, rather than copying directly from the original text. In the case of automatic summarization, abstractive summaries are “more similar to those produced by humans” Giarelis, Mastrokostas, and Karacapilidis (2023) and have become more popular as summarization models have improved.
English datasets for training and evaluating summarization models range from the original Document Understanding Conference (DUC) and Text Analysis Conference (TAC) shared tasks Dang (2006); Dang and Owczarzak (2008), which are relatively small but high quality, to automatically collected datasets, such as CNN/DailyMail Hermann et al. (2015) or XSUM Narayan, Cohen, and Lapata (2018), which rely on collecting news articles paired with their subhead as a proxy for a human-generated summary.
Non-English languages have fewer resources available and most use news articles paired with their subhead Scialom et al. (2020); Varab and Schluter (2021). There are notable exceptions that instead collect human-generated summaries Segarra Soriano et al. (2022); Gonzalez-Agirre et al. (2024), but for most languages there are no high-quality resources available for evaluating automatic summaries.
2.1 Manual Evaluation
Manual evaluation of automatic summaries has long been the gold standard approach, ranging from direct annotation of criteria to the pyramid method Nenkova and Passonneau (2004). This evaluation can be done through directly annotating each summary for certain criteria—coherence of the generated text, relevance of the selected material, etc.—or alternatively by ranking automatic summaries with respect to one of these criteria. In addition, as abstractive summaries have become more common, evaluating their consistency (factuality/faithfulness) has become increasingly important as well Maynez et al. (2020); Vasilyev, Dharnidharka, and Bohannon (2020); Laban et al. (2022).
Importantly, crowdsourcing often leads to poor results when evaluating automatic summarization Gillick and Liu (2010); Hardy, Narayan, and Vlachos (2019); Fabbri et al. (2021) and expert evaluation can also fail to achieve good levels of inter-annotator agreement Steen and Markert (2021); Iskender, Polzehl, and Möller (2021). This can be mitigated to a degree by setting up well-defined annotation guidelines in order to achieve reliable human evaluation of text summarization.
2.2 Automatic Evaluation
Manual evaluation becomes an impossible task as corpus size increases, motivating the need for automatic metrics for summarization evaluation. Reference-based metrics include ROUGE Lin (2004) and its many variants Rankel et al. (2013); Graham (2015), which use n-gram overlaps between reference and automatic summaries. S3 Peyrard, Botschen, and Gurevych (2017) is instead a learned metric that combines the output of other metrics as features and learns to predict human scores directly. Similarly, referenced-based Natural Language Generation (NLG) evaluation metrics, such as BLEU Papineni et al. (2002), BertScore Zhang et al. (2020), CHRF, or CIDEr, are also commonly used.
Reference-free metrics have also been proposed Louis and Nenkova (2009a, b); Saggion et al. (2010). SummaQA Scialom et al. (2019), for example, automatically creates questions from the source document and scores the summary depending on its ability to answer these questions. BLANC Vasilyev, Dharnidharka, and Bohannon (2020) instead measures how much a summary helps a language model predict words from a sentence taken from the original document.
More recently, using LLMs to automatically evaluate summaries—LLM-as-a-Judge—has shown promise Leiter et al. (2023); Kim et al. (2024); Lyu et al. (2025). These models often show higher correlations with human judgments than previous token-based metrics.
However, nearly all of these metrics and models have been designed for English and their ability to correctly evaluate non-English summaries has not been deeply studied. In this line, Lloret, Plaza, and Aker (2018) highlight the importance of not relying solely on summarization evaluation metrics conceived for English texts. It is furthermore nontrivial to transfer many of these metrics to new languages, as they require a number of resources to already exist.
2.3 Meta-evaluation of Automatic Metrics
Meta-evaluation of automatic metrics is necessary to establish the reliability of the metrics. This is typically done by studying the correlation between metric scores and human judgments Lin (2004).
Fabbri et al. (2021), for example, collect human evaluations for 23 English summarization models over four dimensions: coherence, consistency, fluency, and relevance. They then calculate correlations between these annotations and automatic metrics and find that few automatic metrics correlate well with the human annotations.
Despite some studies on Spanish and French Saggion et al. (2010), there has been little meta-evaluation done on other languages. Clark et al. (2023) create a multilingual summary dataset with six languages and also collect binary yes/no crowd-sourced annotations for six criteria, although interannotator agreement (IAA) is quite low, with Krippendorf’s $\alpha$ below 0.5 for 5 of 6 criteria. They then train a referenceless metric and compare it with ROUGE-L, but do not perform a comprehensive study of evaluation metrics. Similarly, Aharoni et al. (2023) create a multilingual summary dataset of 45 languages and crowd-source annotations for two criteria. They then compare ROUGE-L and an NLI-based faithfulness metrics on these data. Finally, Koto, Lau, and Baldwin (2021) meta-evaluate a number of automatic metrics on summaries generated by BERT and pointer-network models in eight languages.
However, none of the meta-evaluations performed on languages other than English study a range of automatic metrics and most of them use automatically extracted subhead-article pairs as a proxy for reference summaries, which limits the reliability of the findings.
3 Methodology
We develop our multilingual summarization evaluation dataset through three main steps. First, we collect news articles from quality sources in both target languages (Section 3.1). Second, we generate a battery of summaries that includes original subheads, human-written summaries, and automatic summaries from various LLMs using different prompt strategies (Section 3.2). Finally, we conduct expert annotation using five criteria through a multi-round process to ensure reliable assessments (Section 3.3).
3.1 Data Collection
Data collection has been conducted with the idea of exploiting some of the most representative data available. We randomly collected news articles covering diverse topics and genres to ensure variety across both languages.
For Basque, we collect 45 articles from Berria, [^2] https://www.berria.eus the main newspaper in this language, including metadata concerning the URL, date of publication, assigned topic category, title of the article, subhead, as well as the body of the text. [^3] The documents were processed with trafilatura Barbaresi (2021). Berria offers its contents under CC BY-SA 4.0 license. [^4] In addition to the 45 articles we annotate, we also collect 22,525 Basque articles from Berria with their subhead which we release with BASSE.
For Spanish, another set of 45 news articles was randomly extracted from the test set of the MLSUM dataset Scialom et al. (2020). The Spanish subset of MLSUM contains news articles of El País newspaper [^5] https://www.elpais.com and provides (news body, subhead, URL) triplets. MLSUM is distributed under MIT license, complying with our needs for redistribution.
3.2 Summary Collection
For each of the 45 articles in both languages, we collected three types of summaries, the final multilingual dataset comprising 90 source articles and 2,040 summaries:
  • Subhead: Following standard practice in news summarization literature Hermann et al. (2015); Narayan, Cohen, and Lapata (2018), we extracted the original subheads from each article to serve as baseline summaries.
  • Human: We collected manually written summaries from our expert annotators. For the first 15 articles, which were used to compute annotation agreement (as explained below in Section 3.3.2), each of the three annotators produced a summary. The remaining 30 articles were divided among the annotators, yielding a total of 75 manual summaries per language.
  • Automatic: We generated automatic summaries using five different LLMs, each prompted in four different ways, resulting in 900 automatic summaries per language. The specific models and prompting strategies are described in the following subsections. This comprehensive approach to summary generation was designed to achieve high variability in summary characteristics, enabling a thorough analysis of summarization quality across different generation methods.
3.2.1 Models
We selected a number of proprietary and open-source LLMs that are able to provide summaries in Basque and Spanish. Proprietary models included Anthropic’s Claude, [^6] https://www.anthropic.com/claude OpenAI’s GPT-4o, [^7] https://openai.com/index/gpt-4o-system-card and Reka AI’s Reka Reka Team et al. (2024). [^8] We used Claude 3.5 Sonnet for Rounds 1, 2, and 3 and Claude 3.5 Haiku for Round 4. Similarly, we used Reka-Core for Rounds 1,2, and 3 and Reka-Flash for Round 4. The open-source options were Cohere’s Command R+ [^9] https://docs.cohere.com/v2/docs/command-r-plus and Meta’s Llama 3.1 70B Instruct Dubey et al. (2024). All models were prompted using their default configurations through their respective native chat interfaces, with the exception of Llama, which was accessed via HuggingFace’s chat interface [^10] https://huggingface.co/chat due to the absence of a proprietary chat platform.
Table 1: Prompts used to collect summaries from LLMs. Note that [doc] is a placeholder for the actual source document, i.e., the original newspaper article.
{tblr}
[…]
3.2.2 Prompts
Along with the selection of those five different models, we experimented with the following four prompts to collect summaries from the LLMs. These prompts, presented below, were designed in order to obtain fairly diverse summaries. Although multilingual LLMs often show stronger performance when prompted in English Shi et al. (2022); Etxaniz et al. (2024a), in early experiments, we found that prompting the LLMs in English led to the summaries also being given in English. Therefore, we opted to prompt the model directly in the language in which it should write the summary. Table 1 shows each prompt in Basque and Spanish, as well as its translation to English and their descriptions are provided below.
  • Base: This prompt asks the LLM to summarize the following document.
  • Core: This prompt asks the model to think about the most important information in the original document and use that information to create a summary for the following document.
  • 5W1H: This prompt, which is a further derivation of Core, uses an approach common in journalism consisting of answering six key questions: the five “W” questions (who, what, when, where, why) and the 1 “H” question (how). This technique “clearly describe[s] key information of news in an explicit manner” Zhang, Chen, and Ma (2019) and allows to capture “the semantic constituents of a sentence which are comparatively simpler to understand and identify” Chakma et al. (2020). The model should then use this information to construct the summary for the following document.
  • Too long, didn’t read (tldr): This prompt appends the common internet approach to summarizing long documents.
3.3 Human Annotation
We followed the annotation procedure of Fabbri et al. (2021), who based their criteria largely on Kryscinski et al. (2019) and Dang (2005). For our experiment, besides the following descriptions of the used criteria, we elaborated a detailed set of annotation guidelines, which we provide in Appendix A. Figure 12 shows real examples of annotated summaries in Basque (examples in Spanish are shown in Appendix B).
Original news article Emakume batek sexu eraso bat jasan zuela salatu zuen atzo goizaldean, Etxebarrin (Bizkaia). Herriko jaiak ospatzen ari dira, eta kalean jarri ohi dituzten komunetan jazo zen erasoa. Etxebarriko Udalak bertan behera utzi zituen atzorako antolatuta zeuden ekitaldi guztiak. Udalak, sanantonioetako jai batzordeak eta txosnek elkarretaratzea egin dute Zintururi plazan, eta ehunka herritar bildu dira ‘Erasorik ez, erantzunik gabe’ lelopean. Ertzaintzak ikerketa abiatu du, eta ez dute inor atxilotu. «Gure jaiak elkartzeko eta dibertitzeko guneak dira, eta gozatu nahi duten pertsona guztiak gonbidatzen ditugu, betiere askatasuna eta bizikidetza errespetatuz. Mezu hau ulertzen ez duten eta bestelako helburuak dituzten pertsonak ez dira ongi etorriak», adierazi du udalak ohar bidez.
Summary: Command R+ with Core prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=red9 Coh & Con Flu Rel 5W1H Avg
2.00 2.67 1.67 2.00 2.00 2.07
Testu honetan, garrantzitsuena da herritarren erantzuna eta errespetua eskatzea. Horregatik, laburtutako testuan, hurrengo frasiak eraginkorrak izan daitezke: “Erasorik ez, erantzunik gabe” lelopea herritarren eragile eta konprometituta dagoena erakusten du. Herriko jai honetan, emakumeen eta pertsonen segurtasuna babesteko, herritar guztiak erantzun bat eman behar dute. Udalak ere, adierazpen baliagarriak egin ditu, ulertu ez duten eta erasoak egiten dituzten pertsonak ez dira ongi etorriak. Horrela, herriko segurtasun eta askatasun boterea eraginkorra sustatzen du. Bertan behera utzi zituen ekitaldi guztiak, udalak erantzukizuna erakusten du, eta herriko jai honetan segurtasun eta askatasun inguruko normak betetzen dituela adierazi nahi du. Horregatik, garrantzitsuena da herritarren erantzuna eta errespetua eskatzea, eta udalak ere, erantzukizuna hartzen duen adierazpen baliagarriak egin du.
Summary: GPT-4o with Base prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=azure9 Coh & Con Flu Rel 5W1H Avg
4.67 4.67 4.67 4.00 5.00 4.60
Hona hemen testua laburtuta: Etxebarrin (Bizkaia), herriko jaietan, emakume batek sexu eraso bat jasan zuen atzo goizaldean kalean jarritako komunetan. Ondorioz, Udalak egun horretarako aurreikusitako ekitaldi guztiak bertan behera utzi zituen. Zintururi plazan elkarretaratzea egin zuten Udalak, jai batzordeak eta txosnek, eta ehunka herritar bildu ziren “Erasorik ez, erantzunik gabe” lelopean. Ertzaintzak ikerketa abiatu du, baina ez da inor atxilotu oraindik. Udalak ohar batean adierazi du jaiak bizikidetza eta askatasuna errespetatuz ospatu behar direla, eta erasoak egiten dituzten pertsonak ez direla ongi etorriak.
Figure 1: Examples of low- and high-quality summaries of a news article in Basque. Scores are the average of three annotators. Source of the article: Berria. [^12] https://www.berria.eus/euskal-herria/ehunka-herritarrek-etxebarriko-sexu-erasoa-salatu-dute_2126343_102.html
3.3.1 Criteria
While maintaining the core criteria of previous work Fabbri et al. (2021), we extended the framework with an additional criterion, 5W1H, which we considered necessary to determine whether the summary contains all relevant information from the source document. The five criteria used in our evaluation are defined as follows:
  • Coherence measures the collective quality of all sentences. The summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to sentence to a coherent body of information about a topic.
  • Consistency refers to the factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Annotators are also asked to penalize summaries that contain hallucinated facts.
  • Fluency scores the quality of individual sentences. Drawing again from the DUC quality guidelines, sentences in the summary “should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read” Dang (2006).
  • Relevance selects the important content from the source. The summary should include only important information from the source document. Annotators are instructed to penalize summaries that contain redundant and excess information.
  • 5W1H measures how well important information is preserved from the source by following a journalistic approach. It evaluates whether key content is presented completely by addressing the six fundamental questions (see prompt 5W1H in Section 3.2.2). The summary should not lack any important information available in the source document. This recall-oriented criterion corresponds roughly to coverage in Koto, Lau, and Baldwin (2020), although it differs in that we calculate 5W1H relative to the original document, whereas they use the reference summary to calculate coverage.
3.3.2 Annotation Procedure
We performed three rounds of expert annotation, where the first two served to refine the annotation guidelines. Given the generally poor agreement of crowdsourced annotations for summarization Gillick and Liu (2010); Fabbri et al. (2021), the annotators were recruited among NLP linguists and engineers who had extensive experience with annotation projects and knowledge of summarization.
In the first round, three annotators evaluated summaries from 10 articles in each language, with each (model, prompt, document) combination being annotated for all five criteria. They additionally annotated the same criteria for the articles’ subheads as baseline summaries and provided their own summaries for each text (which were evaluated by the other two annotators). After calculating agreement scores, annotators discussed the cases of disagreement and refined the guidelines. In the second round, the same set of summaries was re-annotated following the refined guidelines. The third round involved a new set of 5 articles, with all summaries being triple-annotated. Following this round, given the agreement levels achieved, annotators proceeded to individually annotate an additional set of 10 documents each, bringing the total to 45 annotated articles.
After each round of annotation, we calculated interannotator agreement using ordinal Krippendorff’s $\alpha$ Krippendorff (2013) to compare three-way agreement for each of the criteria, and additionally calculated quadratic Cohen’s $\kappa$ Cohen (1960) for pairwise agreement. Results are discussed in the next section.
4 The BASSE Corpus
In what follows, we present BASSE, the manually annotated multilingual corpus of 2,040 news summaries in Basque and Spanish, analyzing interannotator agreement levels, corpus statistics, and qualitative differences across summarization approaches. The diversity of summary generation methods in BASSE enables a robust assessment of summary evaluation metrics under varied conditions.
4.1 Interannotator Agreement
The Krippendorff’s $\alpha$ correlations for each round and language are shown in Table 2. Initial annotations (R#1) yielded moderate agreement levels for Basque across criteria and lower agreement for Spanish. By the final round (R#3), Basque annotations maintained acceptable agreement levels (0.44–0.72), with Spanish remaining at lower values (0.19–0.39). Consistent with prior work Falke et al. (2019); Kryscinski et al. (2020), consistency proved to be the most challenging for annotators to align on—0.44 and 0.19 for Basque and Spanish, respectively, in the final round. This challenge can be attributed to the inherent complexity of evaluating factual alignment, where annotator’s interpretations of source article details sometimes diverge.
While pairwise quadratic Cohen’s $\kappa$ values (Appendix C) follow similar patterns to the Krippendorff’s $\alpha$ scores, raw agreement percentages (Figures 5 and 5) show moderate to high concordance in several categories for Spanish, notably fluency (83–90%) and consistency (65–67%). This apparent contradiction is explained by the score distributions, depicted in Figures 5 and 5: in contrast to Basque summaries, which exhibit a broader distribution across the quality spectrum, Spanish summaries tend to concentrate in the higher range of the quality scale, leading to higher expected agreement and consequently lower chance-corrected scores.
Note that our agreement scores, including those for Spanish annotations, are similar to or exceed those reported in previous large-scale summary evaluation studies of reference Fabbri et al. (2021); Clark et al. (2023); Aharoni et al. (2023).
4.2 Quantitative Analysis
Table 14 presents the statistics of our corpus, which comprises 45 articles with their subheads, 75 human-written summaries, and 180 summaries from each model. The source articles contain 20–40K tokens in total, with Basque showing a markedly lower token count due to its agglutinative nature. A small number of the automatically generated summaries are not written in the target languages Basque and Spanish (6% and 3% respectively). Summary vocabulary sizes range between 3K and 10K tokens, and they vary considerably in length and richness across different approaches.
We analyze the diversity of summarization strategies through two key metrics. Compression ratio indicates summary conciseness with respect to the source article Segarra Soriano et al. (2022), while novel 2-grams measure the percentage of new word pairs in the summary Kryściński et al. (2018).
As shown in Figure 6, subheads achieve the highest compression (10–100$\times$), while human and model summaries typically compress around 6$\times$, with some notable exceptions. GPT-4o has occasionally produced expansive content matching or exceeding source length, while Llama has consistently generated more compressed summaries.
The novel 2-gram analysis shows that model-generated summaries are more abstractive than human-written ones, particularly from Claude (though it must be noted that this finding is partially influenced by Claude’s outputs in languages other than the target one). In contrast, GPT-4o has sometimes opted for a more extractive approach, closely reproducing the original text as mentioned earlier.
Table 2: Ordinal Krippendorff’s $\alpha$ across triple-annotated rounds (R#). Rounds 1 and 2 were conducted on the same set of summaries (Doc), before and after guideline discussion respectively, while round 3 involved a new set of summaries.
{tblr}
[…]
2 210 0.59 0.63 0.75 0.54 0.64 0.66 0.37 0.35 0.49 0.58
3 105 0.66 0.44 0.70 0.63 0.72 0.29 0.19 0.34 0.20 0.39
[…]
Figure 2: Agreement percentage in Basque R#3
Figure 3: Agreement percentage in Spanish R#3
Figure 4: Score distribution across annotators in Basque R#3
Figure 5: Score distribution across in Spanish R#3
Table 3: Statistics of the BASSE corpus, including the source articles and their summaries. We report the number of documents (Doc)—and, of those, how many are not in the target language (¬ eu or ¬ es)—, total tokens (Tok), vocabulary size (Voc), and sentences per document (s/d). Splitting and tokenization were done with stanza Qi et al. (2020) and language identification with lingua-py. [^14] https://github.com/pemistahl/lingua-py
{tblr}
[…]
Human summary 75 0 8,867 3,428 6 75 0 11,635 3,180 5
LLM summary 900 54 133,755 15,755 10 900 27 158,945 10,844 8
by model
  Claude 180 38 29,602 7,589 12 180 26 32,301 6,226 10
  Cmd R+ 180 16 37,236 7,945 13 180 1 30,460 4,724 7
  GPT-4o 180 0 30,859 7,235 12 180 0 29,647 4,962 7
  Llama 180 0 14,671 3,279 6 180 0 33,918 4,401 8
  Reka 180 0 21,387 4,658 7 180 0 32,619 4,858 7
by prompt
  Base 225 22 43,452 9,965 12 225 0 42,242 5,836 8
  Core 225 0 24,746 5,378 7 225 0 36,059 5,349 7
  5W1H 225 0 37,273 6,740 15 225 0 49,885 5,625 10
  tldr 225 32 28,284 7,483 7 225 27 30,759 6,099 6
[…]
Figure 6: Distribution of compression ratios (x axis) and novel 2-grams (y axis) for different summarization approaches, in Basque (eu) and Spanish (es). Higher compression values ($\rightarrow$) indicate more concise summaries, while higher novel 2-gram values ($\uparrow$) suggest more abstractive summarization.
4.3 Qualitative Analysis
Table 4: Mean human evaluation scores on a 5-point Likert scale for summaries across the five evaluation criteria, and their absolute average (Avg). Highest scores for each criterion and computing class are highlighted in bold.
{tblr}
[…]
Human 4.77 4.94 4.85 4.51 4.69 4.75 4.91 4.88 4.80 4.48 4.34 4.68
LLM 3.46 4.29 3.87 3.74 4.06 3.89 3.88 4.68 4.82 4.14 4.18 4.34
by model
  Claude 3.17 4.61 3.83 3.54 4.57 3.95 3.17 4.71 4.40 3.81 4.40 4.10
  GPT-4o 3.88 4.52 4.55 3.78 4.62 4.27 4.13 4.77 4.94 4.37 4.37 4.52
  Reka 3.72 4.20 4.05 4.02 4.14 4.03 4.21 4.49 4.87 4.26 4.17 4.40
  LLama 3.1 3.59 4.50 4.29 4.31 3.45 4.03 3.78 4.79 4.93 3.98 3.94 4.29
  Cmd R+ 2.95 3.65 2.64 3.07 3.53 3.17 4.10 4.66 4.94 4.29 4.02 4.40
by prompt
  Base 3.78 4.20 3.70 3.36 4.19 3.85 4.22 4.67 4.91 4.12 4.20 4.42
  Core 3.61 4.33 4.05 3.73 3.64 3.87 4.11 4.73 4.93 4.08 3.94 4.36
  5W1H 2.62 4.22 4.10 3.64 4.39 3.79 3.03 4.58 4.96 4.03 4.65 4.25
  tldr 3.82 4.42 3.64 4.24 4.01 4.03 4.14 4.75 4.47 4.35 3.94 4.33
[…]
The manual evaluation of the summaries (see Table 4; detailed results per model and prompt combination are available in Appendix D) shows significant quality differences across summarization approaches. While human-written summaries achieve the highest average scores (4.75 for Basque, 4.68 for Spanish), that is not the case for the subheads, which fail at covering key information (criterion 5W1H).
Among the model-generated summaries, we observe a clear linguistic divide, with Spanish summaries generally achieving higher fluency scores. This gap is most pronounced for Command R+, where poor fluency in Basque (2.64 compared to 4.92 in Spanish) consequently impacts its consistency. GPT-4o shows the best performance in both languages, although all models show varying performance patterns across different prompting strategies.
Most LLMs achieve consistency scores comparable to humans, and some models occasionally surpass human performance in relevance and 5W1H coverage. In fact, the 5W1H prompt seems to compel models to provide a better coverage of key information.
All models normally generate summaries with a suitable format. However, Claude shows a tendency to generate bullet-point summaries, which negatively impacts its coherence scores, and occasionally produces English text with the tldr prompt, as noted earlier. Additionally, it is the most prone to generating introductory sentences for the summaries, e.g., “Aquí está el resumen del texto:” (Here is the summary of the text:), which negatively impacts its relevance scores.
5 Evaluation
In this section, we calculate the correlation between the human annotations from Section 4 and a series of evaluation metrics and LLM-as-a-Judge commonly used for evaluating NLG. For rounds with more than one annotation per summary, we take the average of the three values for each criterion. We calculate both Spearman’s ($\rho$) and Kendall’s ($\tau$) rank correlation coefficients at system-level following Louis and Nenkova (2013) and Fabbri et al. (2021), but report only Spearman in the subsequent section, as the results do not vary depending on the coefficient.
5.1 Evaluation Metrics
We select a subset of metrics used in SummEval Fabbri et al. (2021) which require minimal changes to work for languages other than English. Due to their dependence on specialized English training data or English-only models, several of the original metrics are excluded, namely, BLANC Vasilyev, Dharnidharka, and Bohannon (2020), S3 Peyrard, Botschen, and Gurevych (2017), SummaQA Scialom et al. (2019), MoverScore Zhao et al. (2019), Sentence Mover’s Similarity Clark, Celikyilmaz, and Smith (2019), and SUPERT Gao, Zhao, and Eger (2020). For several other metrics (i.a., ROUGE-we, BertScore, METEOR, CIDEr, and Data Statistics, explained below), we implement the necessary changes to allow for evaluation in Spanish or Basque. Before passing the candidate or reference summaries to each metric, we perform metric specific preprocessing of the texts (see details in the description of each metric).
  • ROUGE Lin (2004) is a series of metrics that calculate token subsequence overlap between a candidate summary and a set of reference summaries. We calculate ROUGE-{1–4}, ROUGE-L, and ROUGE-su* variants.
  • ROUGE-we Ng and Abrecht (2015) extends the original ROUGE metric by incorporating soft matching based on cosine similarity of word embeddings. We use FastText embeddings for both Basque and Spanish. [^15] https://fasttext.cc/docs/en/crawl-vectors
  • BertScore Zhang et al. (2020) calculates the similarity of two texts as the average token-level cosine similarities between greedily aligned tokens in each text. For our experiments, we replace BERT with multilingual BERT base uncased [^16] https://huggingface.co/google-bert/bert-base-multilingual-uncased and report precision (p), recall (r), and F1 (f) scores.
  • BLEU Papineni et al. (2002) is a corpus-level metric originally designed for machine translation evaluation. It calculates average n-gram precision, coupled with a brevity penalty. We specifically use the SacreBLEU implementation Post (2018).
  • CHRF Popović (2017) calculates character-level n-gram overlaps between the candidate summary and the reference summary.
  • METEOR Banerjee and Lavie (2005) computes an alignment between candidate and reference sentences, mapping unigrams in the generated summary to 0 or 1 unigrams in the reference, based on stems, synonyms, and paraphrase matches. Precision and recall are computed and reported as a harmonic mean. We compute Multilingual METEOR (mMETEOR) by setting the language to other and only lower-casing the input.
  • CIDEr Vedantam, Zitnick, and Parikh (2015) is a corpus-level metric originally designed for image description evaluation. It calculates the average cosine similarity between TD-IDF weighted n-gram representations of the candidate summary and a set of reference summaries, taking the average score of 1–4 grams. As this metric applies stemming, we use the Basque and Spanish versions of the Snowball Stemmer Porter (2001).
  • Data Statistics Grusky, Naaman, and Artzi (2018) calculates a series of metrics originally designed to describe how extractive a dataset is. This includes i) Coverage, which measures how derivative a summary is; ii) Density, which measures how well a summary can be described as a series of extractions from the original text or iii) Compression, which is the ratio of the length of the summary divided by the length of the original text. SummEval further introduces n-gram related novelty—novel {1–3} gram—and redundancy scores—repeated {1–3} gram. Finally, we also include the length of the generated summaries.
5.2 LLM-as-a-Judge
We further experiment with using LLM-as-a-Judge systems to predict the five criteria we use for human evaluation. Currently, no LLM-judge models officially support Basque or Spanish evaluation specifically. Therefore, we selected a range of models that have demonstrated strong performance as judges in English and that offer some multilingual capabilities. Our selection represents diverse characteristics: open-source versus proprietary models, and varying parameter sizes and specializations. Following these selection principles, we evaluate these seven models as judges:
  • Prometheus 2 Kim et al. (2024) is an open-source LLM trained to evaluate the output of other models following an annotation rubric. Although Prometheus 2 is primarily trained for English, its backbone model Mistral-7B-Instruct-v0.2171717https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 is a multilingual model able to process Spanish, but not Basque. We experiment with both the 7B and 8x7B parameter models.
  • Selene Mini Alexandru et al. (2025) is similarly an open-source LLM [^18] https://huggingface.co/AtlaAI/Selene-1-Mini-Llama-3.1-8B that uses a mix of reinforcement learning and supervised fine-tuning to train Llama3.1-8B-Instruct Dubey et al. (2024) to act as a general-purpose evaluator. For English, this model shows better performance than previous open-source models and is comparable to small closed-source models.
  • Qwen2.5 Qwen et al. (2025) is a set of open instruction-tuned LLMs that have shown promise as judge models in English Gu et al. (2025). It officially supports Spanish, and authors report transfer to low-resource languages through translated data. We specifically use the 7B and 72B instruction-tuned models.
  • GPT-4o is a general-purpose commercial LLM that has shown strong performance as a judge model in English Huang et al. (2024). It has also shown robust performance in other Spanish and Basque tasks Achiam et al. (2023); Etxaniz et al. (2024a, b). We experiment with GPT-4o mini and GPT-4o. [^19] The specific models used were gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06.
We use the codebase from Prometheus-Eval [^20] https://github.com/prometheus-eval/prometheus-eval and format the prompt to include the original instruction, the annotation rubric for the criteria, a human-generated reference answer, and the summary to be evaluated. We evaluate only one criterion at a time. Although the original documents and summaries are in either Basque or Spanish, we use English for the instructions and annotation rubric, as preliminary results indicated that several models struggle to follow the directions otherwise. The full prompt template used for all judge models is shown in Figure 7.
You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
###Task Description:
An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given.
1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general.
2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric.
3. The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)"
4. Please do not generate any other opening, closing, and explanations.
###The instruction to evaluate:
{instruction}
###Response to evaluate:
{response}
###Reference Answer (Score 5):
{reference answer}
###Score Rubrics:
{rubric}
###Feedback:
Figure 7: Prompt from Prometheus-Eval used for LLM-as-a-Judge evaluation. Note that instruction, response, reference answer, and rubric are placeholders that are filled with specific information for each example.
6 Results
In this section, we answer the research questions by analyzing Spearman’s $\rho$ correlations between model rankings derived from automatic evaluations and those from human judgments. Selected results are shown in Figure 8, while full results of both Spearman’s $\rho$ and Kendall’s $\tau$ rank correlation coefficients can be consulted in Appendix E.
RQ1: Do automatic metrics for evaluating summarization correlate well with human judgments in Basque and Spanish?
Overall, we observe notable patterns across languages in metric performance. Traditional metrics show similar trends in both Basque and Spanish for coherence, consistency, and relevance, with several metrics showing strong correlations with human judgments, despite the gaps in annotator agreement for these criteria. Notably, fluency and 5W1H—the criteria with highest human agreement in both languages—result in nearly opposite tendencies between languages, with better correlations in Basque. On the other hand, LLM-as-a-Judge models form different patterns: they follow similar trends in coherence, relevance and 5W1H, and almost inverse trends in consistency and fluency. Looking at each criterion in detail:
Figure 8: Spearman’s model-level rank correlation between automatic metrics and human annotations, in Basque (eu) and Spanish (es).
Automatic metrics have a significant correlation with coherence with overall absolute mean Spearman correlations of 0.548 and 0.454. BertScore-p gives good results on both Basque and Spanish (0.746/0.774), while the best performing metrics differ by language: BertScore-f for Basque (0.814) and CIDEr for Spanish (0.786). Although ROUGE-L has a generally high correlation in both languages (0.738/0.675), other more common variants of ROUGE (1–4) have lower correlations. Finally, in Spanish, mROUGE-we gives the third best result (0.752). Interestingly, for this criterion, the automatic metrics perform on par with proprietary LLM-as-a-Judge models.
Consistency has the lowest absolute mean correlations between metrics and human judgments (0.201/0.257), indicating that the metrics struggle to capture this criterion. In Basque, GPT-4o models achieve the best correlations (0.574 and 0.509) after which follows bigram repetition (-0.455). In Spanish, novel bigrams and trigrams give the strongest absolute correlations (-0.553/-0.617) and no LLM-as-a-Judge achieves more than 0.105.
Fluency similarly has low overall absolute mean correlations (0.367/0.164). In Basque, however, GPT-4o achieves 0.761 and ROUGE-{1–3} have high correlations (0.556, 0.677, and 0.606, respectively). In contrast, these same metrics have poor correlations in Spanish (-0.055, -0.280, -0.071, and 0.100). The best-performing metric for that language is CIDEr (-0.429). The same metric often gives opposite results in Basque and Spanish, as is the case of most ROUGE variants or BertScore.
Relevance has relatively high absolute mean correlations (0.501/0.358). Statistics-based metrics have generally high correlations in both languages—e.g., length (-0.887/-0.621), novel unigram (-0.780/-0.686), novel bigram (-0.756/-0.484)—, while BertScore-p (0.737/0.801) and CIDEr (0.677/0.654) also capture this criterion well across both languages. Most ROUGE variants, however, perform poorly, as do all judge models.
Finally, 5W1H has moderate absolute mean correlations (0.375/0.463). LLM-as-a-Judge models capture this criterion well, with $\rho$ values above 0.5 for all models except Prometheus2 7B on Spanish. Compression is the single automatic metric that best captures 5W1H across both languages (-0.665/-0.792), although coverage (-0.848) and novel unigram (0.816) perform the best in Spanish.
In summary, several metrics correlate strongly for both languages ($\rho>$ 0.6): For coherence and relevance, BertScore-p shows a strong correlation in both languages, and compression performs well for 5W1H. In Basque, ROUGE-2 and ROUGE-3 capture fluency well, while no automatic metric performs well on fluency in Spanish, likely due to the low variance in scores. Similarly, there is no metric that consistently captures consistency across both languages. Metrics originally designed for machine translation evaluation (namely, BLEU, CHRF, and METEOR) do not stand out for any specific summarization criterion, generally exhibiting weak correlations across most evaluation dimensions in both languages.
RQ2: Do LLM-as-a-judge models have the ability to replicate human judgments for summarization?
In general, proprietary LLM-as-a-Judge models (GPT-4o Mini and GPT-4o) give the highest correlations with human judgments on most criteria. GPT-4o, for example, gives the highest correlation in four of five criteria for Basque, with $\rho$ ranging from 0.512 to 0.909 and one of five in Spanish, with $\rho$ between -0.055 and 0.876. The smaller GPT-4o mini similarly performs well, ranging from -0.097 to 0.874 in Basque and -0.28 to 0.783 in Spanish.
The open-source LLM-as-a-Judge models, on the other hand, do not generally outperform the automatic metrics, with the exception of Selene Mini with $\rho$ of 0.812 for 5W1H in Basque. Both Prometheus 2 8x7B and Qwen2.5-72B-Instruct perform as poorly as their smaller versions, suggesting that the low correlation achieved is not due to model size. Furthermore, both Prometheus 2 judges show better—though still very weak—correlations in Basque than in Spanish, which is particularly notable given that the backbone model of Prometheus, Mistral, officially supports Spanish but not Basque.
Summing up, proprietary models show relatively strong correlations with human judgments (GPT-4o above all, and GPT-4o mini to a lesser degree) while open-source models currently are only able to capture 5W1H.
RQ3: Which summary criteria (coherence, consistency, fluency, relevance, 5W1H) do state-of-the-art LLMs struggle with in Basque and Spanish?
As noted earlier, our goal was not to produce optimal summaries but rather to gather a diverse range of summary qualities for meta-evaluation purposes. We deliberately avoided hyperparameter tuning or extensive prompt engineering that may have yielded better results. Thus, the following results better represent what users can expect out of the box from these models.
In Basque, coherence is the criterion with the lowest score, with an average of 3.46/5.00 over all model prompt combinations, followed by relevance (3.74), while consistency (4.29) and 5W1H (4.06) score slightly better. That is, Basque summaries are fairly faithful to source content and comprehensive in their information coverage, yet they struggle with logical flow and content relevance. The moderate fluency average score (3.87) masks considerable variation in linguistic competence between different models.
In Spanish, the models generally achieve better scores. Fluency has the highest score (4.82), followed by consistency (4.68). Coherence is lowest (3.88) while relevance (4.14) and 5W1H (4.18) lie in the middle. In other words, Spanish summaries excel in linguistic quality and faithfulness while providing mostly complete and relevant information, although logical connectivity between ideas needs some improvement.
Overall, for the models tested, coherence is the most problematic criterion currently in both languages, while fluency and relevance may be challenging specifically in Basque. Detailed results are available in Appendix D.
7 Discussion
We next examine the limitations inherent in our methodology, and analyze the contributions of our work, discussing the implications of the performance observed across different metrics and evaluation methods.
7.1 Methodological Considerations
Although our corpus collection and annotation methodology are the same for both Basque and Spanish, the outcome differs between languages. In Basque, the summaries we collected are generally of poorer quality, but the IAA is high, while for Spanish the quality of all summaries is quite good, but the IAA is lower. This seems to be an artifact of the higher quality of the summaries in Spanish. Annotators rarely used the lower end of the scale (1–3) for several criteria, and the expected agreement is therefore higher, affecting IAA scores, despite raw agreement often being over 80%.
Beyond language-specific differences, the domain of our study offers important context for interpreting our findings. This meta-evaluation is performed on newswire data, and our results may not necessarily transfer well to other domains or formulations of summarization, as newswire texts tend to be written in a declarative style, avoid negative constructions, and explicitly represent 5W1H information. However, the standard coherence, consistency, fluency and relevance metrics have been successfully used to annotate summaries in other domains, and it is likely that the new 5W1H metric can be easily adapted to different types of texts according to the relevance of each of the 5W1H elements in text.
Finally, note that most of the news articles in our dataset have a single human-generated summary, which limits our ability to exploit the variability of human summaries. However, given a set annotation budget and the goal of creating a large and varied set of summaries, previous research has shown it is preferable to increase the number of instances, even if they are single annotated Steen and Markert (2021).
7.2 On Contributions and Findings
A key proposal of our research is the introduction of the 5W1H evaluation criterion, which has shown promise as a recall-oriented measure of important information from the original document. One may expect it to be the inverse of relevance—a precision-oriented measure of information transfer—but our experiments suggest that there is no direct correlation between them ($\rho$ = -0.16 and -0.34 at model level) and different metrics capture each criterion better. In fact, while LLM-as-a-Judge models generally have $\rho>$ 0.561 for 5W1H, they have close to random correlations for relevance.
Building on the evaluation criteria, we observed that proprietary LLM-as-a-Judge models generally outperform traditional metrics, especially on coherence and 5W1H, while the open-source LLM-as-a-Judge models perform much worse. This performance gap is likely due to the fact that neither Prometheus 2 nor Selene Mini are trained as judges for languages other than English, while the proprietary models may have such training. Our experiments thus serve as an interesting testing ground for judge LLMs in cross-lingual contexts. Although multilingual models often can transfer knowledge across languages, this capability does not appear to extend to summary evaluation in our study. This does not indicate that multilingual models are not possible, but rather points to the importance of developing truly multilingual judge models to enable high quality evaluation in non-English languages.
Finally, we observed that—despite LLMs being capable of producing mostly faithful summaries—consistency is currently poorly captured by LLM-as-a-Judge models, as well as by all automatic metrics, contrary to previous findings in English Fabbri et al. (2021). This is likely due to the choice of summarization models, as previous meta-evaluations have included summaries from extractive summarization models and simpler abstractive models. Current abstractive summarization models often have subtle inconsistencies Kryscinski et al. (2020); Yang et al. (2024), potentially making them more difficult to capture.
8 Conclusion
In this work, we present BASSE, a dataset of 2,040 automatic summaries in Basque and Spanish that we collected and annotated manually across five criteria on a 5-point Likert scale: coherence, consistency, fluency, relevance and 5W1H. The interannotator agreement study performed on a subset of triple-annotated summaries yields better scores than those found in prior large-scale summarization evaluation research. We used this dataset to perform a meta-evaluation of traditional automatic metrics and LLM-as-a-Judge models beyond English contexts. Further, we share the first automatically collected summarization dataset in Basque, which comprises 22,525 article-subhead pairs.
We observed a hierarchy of evaluation approaches: proprietary LLM-as-Judge models demonstrate the strongest overall correlations with human judgments, particularly for coherence and 5W1H coverage. On the other hand, traditional metrics like BertScore-p, ROUGE variants, and compression ratio show strong criterion-specific performance but lack consistency across all dimensions. At the bottom of this hierarchy, open-source judge LLMs perform notably poorly, suggesting that current multilingual capabilities do not effectively transfer to the task of summarization evaluation in non-English languages.
Among the evaluation criteria, coherence and relevance were most reliably captured by automatic metrics, with BertScore-p showing strong correlations across both languages. The 5W1H criterion was consistently assessed by proprietary judge models. Fluency exhibited language-dependent results, with several metrics performing well for Basque but struggling with Spanish. Consistency proved the most challenging criterion overall, with no metric or model showing strong correlations across both languages, reflecting the difficulty of detecting subtle factual misalignments in today’s abstractive summarization systems’ output.
We release the dataset and code with the hope that it will be useful for developing further multilingual evaluation methods for automatic summarization.
[…]
Annotation Guidelines
Coherence
Coherence refers to the collective quality of all sentences. We align this dimension with the DUC quality question Dang (2005) of structure and coherence whereby ‘the summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentence to sentence to a coherent body of information about a topic.’ To clarify, does the summary have a coherent flow of ideas and does it have connectors to explicitly define the relationships between ideas? It does not matter if there are grammatical errors or incorrect information. If the summary is simply a list of events, annotators should penalize it. Annotators do not penalize single sentences for not having connectors.
Coherence is rated 1-5 according to the following guidelines:
  1. 1.
    The summary contains a bullet point list of events and there is no internal consistency (e.g., random words).
  2. 2.
    The summary contains a bullet point list of events, but there is no consistency between the phrases in each bullet point (e.g., 1. Keyboard, 2. Watching TV).
  3. 3.
    The summary contains a bullet point list of events, but the phrases in each bullet point are well developed.
  4. 4.
    The summary contains implicit coherence, but not explicit. That means that there are no temporal connectors or discourse markers, but the sentences have a temporal/causal flow.
  5. 5.
    The summary contains explicit coherence via textual resources, such as temporal connectors or discourse markers that define the relationship between ideas and sentences.
Consistency
Consistency refers to the factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Annotators should also penalize summaries that contain hallucinated facts. If the summary contains information not found in the original document, annotators should penalize it. For temporal expressions (today, yesterday, this year), if the expression is consistent with the original information, annotators should assume that the summary is consistent and should not penalize.
Consistency is rated 1-5 according to the following guidelines:
  1. 1.
    The summary does not contain any ideas from the original text.
  2. 2.
    The summary contains a large amount of incorrect information.
  3. 3.
    The summary contains several incorrect pieces of information.
  4. 4.
    The summary contains one incorrect piece of information.
  5. 5.
    The summary is completely factual.
Fluency
Fluency refers to the quality of individual sentences. Drawing again from the DUC quality guidelines, sentences in the summary ‘should have no formatting problems, capitalization errors or obviously ungrammatical sentences (e.g., fragments, missing components) that make the text difficult to read.’ Is the summary well written following the grammar and spelling conventions for the language? If it has grammatical or spelling errors, annotators should penalize it. However, this criteria only takes grammar and fluency into account. While the style can sometimes be forced, if it is grammatically correct, we give full points.
Fluency is rated 1-5 according to the following guidelines:
  1. 1.
    There are many important grammatical errors or the summary mixes language varieties/dialects in a very unnatural and uncomfortable way or the summary is not written in the language it is asked for.
  2. 2.
    There are various important grammatical errors or the summary mixes language varieties/dialects in an unnatural and uncomfortable way.
  3. 3.
    There are grammatical or spelling errors of lesser importance or the summary mixes language varieties/dialects in a somewhat unnatural and uncomfortable way.
  4. 4.
    There are few grammatical or spelling errors.
  5. 5.
    There is no grammatical or spelling errors and the variety of language/dialect used is coherent.
Relevance
Relevance refers to the selection of important content from the source. The summary should include only important information from the source document. Annotators should penalize summaries that contain redundancies and excess information. They should also penalize metalinguistic phrases generated by the model, such as ‘Here is the summary:’ (-1). Finally, they should also penalize the presence of subjective opinions if they are not part of the original text (-1).
Relevance is rated 1-5 according to the following guidelines:
  1. 1.
    The summary is the original text or only contains irrelevant information.
  2. 2.
    The summary contains a large amount of irrelevant information.
  3. 3.
    The summary contains a mix of relevant and irrelevant information.
  4. 4.
    The summary contains mostly relevant information.
  5. 5.
    All information in the summary is relevant.
5W1H
5W1H refers to the maintenance of all important information (the Ws in 5W1H - who, what, when, where, why, how) from the source document. The summary should not lack any important information available in the source.
5W1H is rated 1-5 according to the following guidelines:
  1. 1.
    The summary contains no relevant W from the original text.
  2. 2.
    The summary contains only 1-2 Ws.
  3. 3.
    The summary lacks several relevant Ws.
  4. 4.
    The summary is lacking one relevant W.
  5. 5.
    The summary contains all the relevant Ws from the original text.
[…]
Annotation Examples
Figures 22 provides an example of low- and high-quality summaries in Spanish, along with their corresponding average scores across all five evaluation criteria.
Original news article La Policía ha detenido en Alemania al joven de 17 años que estaba siendo buscado como presunto autor de la muerte de su hermana , de 14 , cuyo cadáver fue hallado en el domicilio familiar de Mataró , Barcelona , el pasado junio . El joven ha sido trasladado ya a España y ha sido puesto a disposición de la fiscalía de menores , según fuentes cercanas al caso . El menor estaba siendo buscado desde el pasado mes de junio , después de que se encontrara en el domicilio familiar el cadáver de su hermana , de 14 años , con heridas de arma blanca . El crimen ocurrió cuando la madre de ambos menores , de origen ruso , se encontraba de viaje , por lo que los dos adolescentes se habían quedado solos en la vivienda , situada en el barrio Cerdanyola de Mataró . Alertada por no poder contactar con sus hijos , la mujer pidió a unos amigos que acudieran a su domicilio , donde estos encontraron el cadáver de la menor , sin rastro de su hermano mayor.
Summary: Llama 3.1 70B Instruct with 5W1H prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=red9 Coh & Con Flu Rel 5W1H Avg
3.00 3.67 5.00 3.67 4.67 4.00
Aquí está el resumen del texto usando el método de las 5W1H: Qué: La Policía ha detenido al joven de 17 años como presunto autor de la muerte de su hermana de 14 años, cuyo cadáver fue hallado en su domicilio familiar en Mataró, Barcelona, con heridas de arma blanca. Quién: El joven de 17 años, cuya identidad no ha sido revelada debido a su condición de menor de edad. Cuándo: El crimen ocurrió en junio pasado, y el joven fue detenido recientemente y trasladado a España para ser puesto a disposición de la fiscalía de menores. Dónde: En el domicilio familiar situado en el barrio Cerdanyola de Mataró, Barcelona. Por qué: Las razones detrás del crimen aún no están claras y son objeto de investigación. Cómo: El joven supuestamente atacó a su hermana con un arma blanca, causándole heridas que resultaron fatales. Este resumen proporciona una visión general concisa del incidente utilizando las preguntas básicas de las 5W1H para transmitir la información clave.
Summary: GPT-4o with Base prompt {tblr} width=colspec=*6X,row1=c,font=,row2=c,font=,rowsep=0pt,colsep=0pt,column6=bg=azure9 Coh & Con Flu Rel 5W1H Avg
4.33 5.00 5.00 4.67 5.00 4.80
La Policía ha detenido en Alemania a un joven de 17 años, sospechoso de haber matado a su hermana de 14 años en Mataró, Barcelona, el pasado junio. El joven ya ha sido trasladado a España y puesto a disposición de la fiscalía de menores. El crimen ocurrió cuando su madre estaba de viaje y los adolescentes se quedaron solos en su casa. La madre, al no poder contactar con ellos, pidió a unos amigos que fueran a su domicilio, donde encontraron el cadáver de la menor con heridas de arma blanca y sin rastro del hermano.
Figure 9: Examples of low- and high-quality summaries of a news article in Spanish. Scores are averages of three annotators. Source of the article: El País. [^22] http://elpais.com/ccaa/2019/09/28/catalunya/1569665675_576416.html
[…]
Interannotator Agreement
Figures 11 and 11 show interannotator agreements in terms of quadratic Cohen’s $\kappa$ for the final annotation round (Round #3) in Basque and Spanish respectively.
Figure 10: Quadratic Cohen’s $\kappa$ in Basque R#3
Figure 11: Quadratic Cohen’s $\kappa$ in Spanish R#3
[…]
Summarization Evaluation
Table 5 presents the average human evaluation scores for summaries on a 5-point Likert scale across the five evaluation criteria: coherence (Coh), consistency (Con), fluency (Flu), relevance (Rel), and 5W1H coverage. The table reports scores for the original article subheads, human-written summaries, and the summaries generated by five different LLMs with four prompting strategies.
Table 5: Mean human evaluation scores on a 5-point Likert scale for summaries across the five evaluation criteria, and their absolute average (Avg). Highest scores for each criterion and computing class are highlighted in bold.
{tblr}
[…]
 Base 3.20 4.71 3.13 3.04 4.60 3.74 3.43 4.73 4.96 3.86 4.34 4.26
 Core 3.73 4.57 4.47 3.76 4.30 4.17 3.59 4.81 4.96 3.81 4.21 4.28
 5W1H 2.51 4.36 4.61 3.45 4.85 3.96 2.67 4.59 4.98 3.80 4.76 4.16
 tldr 3.24 4.80 3.10 3.92 4.53 3.92 3.00 4.70 2.69 3.78 4.30 3.69
GPT-4o
 Base 4.41 4.57 4.59 2.61 4.83 4.20 4.53 4.78 4.96 4.34 4.33 4.59
 Core 4.19 4.50 4.56 4.36 4.27 4.37 4.48 4.80 4.93 4.39 4.30 4.58
 5W1H 2.82 4.41 4.56 3.56 4.81 4.03 3.05 4.71 4.93 4.18 4.65 4.31
 tldr 4.08 4.59 4.47 4.60 4.58 4.46 4.47 4.79 4.94 4.57 4.21 4.60
Reka
 Base 4.12 4.10 3.96 3.91 4.40 4.10 4.64 4.39 4.83 4.26 4.16 4.46
 Core 3.88 4.10 4.12 3.73 3.85 3.94 4.48 4.63 4.84 4.13 4.03 4.42
 5W1H 2.81 4.19 4.13 3.87 4.32 3.86 3.30 4.33 4.93 4.10 4.63 4.25
 tldr 4.07 4.39 3.99 4.57 3.97 4.20 4.41 4.61 4.87 4.54 3.87 4.46
Llama 3.1
 Base 3.94 4.33 4.31 4.47 3.39 4.09 4.07 4.82 4.87 3.95 4.09 4.36
 Core 3.67 4.61 4.27 4.02 3.01 3.92 3.53 4.79 4.93 3.61 3.73 4.12
 5W1H 2.64 4.49 4.23 3.83 4.09 3.86 3.09 4.75 4.95 3.87 4.54 4.24
 tldr 4.09 4.56 4.36 4.90 3.30 4.24 4.42 4.81 4.98 4.50 3.41 4.43
Cmd R+
 Base 3.25 3.31 2.50 2.79 3.75 3.12 4.44 4.61 4.92 4.17 4.07 4.44
 Core 2.56 3.88 2.82 2.78 2.79 2.97 4.49 4.64 4.97 4.44 3.45 4.40
 5W1H 2.33 3.64 2.95 3.50 3.88 3.26 3.06 4.53 5.00 4.18 4.67 4.29
 tldr 3.64 3.78 2.30 3.21 3.69 3.32 4.42 4.84 4.86 4.36 3.90 4.48
[…]
Metric Correlation
Tables 6 and 7 present correlation values between automatic evaluation metrics and human annotations for Basque and Spanish respectively. For each automatic metric and human annotation criterion, both Kendall’s $\tau$ and Spearman’s $\rho$ rank correlation coefficients are reported. Rank correlation coefficients have been computed at model level.
Table 6: Rank correlations between automatic metrics and human annotations for the Basque data. Largest three correlations for each criteria in bold.
{tblr}
[…]
ROUGE-2 0.453 0.216 0.501 0.484 0.095 0.630 0.223 0.677 0.534 0.069
ROUGE-3 0.332 0.201 0.455 0.364 0.005 0.463 0.194 0.606 0.474 -0.042
ROUGE-4 0.295 0.164 0.364 0.305 -0.105 0.409 0.148 0.487 0.421 -0.179
ROUGE-L 0.537 0.111 0.311 0.526 0.032 0.738 0.126 0.496 0.683 0.015
ROUGE-su* 0.505 0.121 0.343 0.453 0.232 0.695 0.181 0.557 0.552 0.271
mROUGE-we 0.442 0.111 0.290 0.411 0.000 0.689 0.088 0.451 0.523 -0.027
[…]
BertScore-r 0.316 0.016 0.343 0.053 0.421 0.441 -0.071 0.540 0.045 0.525
BertScore-f 0.611 0.079 0.322 0.432 0.042 0.814 0.097 0.475 0.543 0.033
[…]
CHRF -0.147 0.132 0.090 -0.011 0.400 -0.269 0.190 0.108 -0.032 0.523
mMETEOR 0.347 0.026 0.332 0.211 0.368 0.462 -0.005 0.527 0.208 0.406
CIDEr 0.211 0.047 0.195 0.495 -0.084 0.302 0.079 0.297 0.677 -0.108
[…]
Novel 1-gram -0.326 -0.142 -0.206 -0.611 0.305 -0.537 -0.117 -0.351 -0.780 0.412
Novel 2-gram -0.368 -0.142 -0.206 -0.568 0.305 -0.591 -0.127 -0.329 -0.756 0.438
Novel 3-gram -0.421 -0.121 -0.195 -0.474 0.295 -0.630 -0.104 -0.297 -0.668 0.462
Rep. 1-gram -0.474 -0.280 -0.332 -0.421 -0.242 -0.684 -0.416 -0.497 -0.600 -0.349
Rep. 2-gram -0.411 -0.343 -0.185 -0.463 -0.179 -0.597 -0.455 -0.282 -0.614 -0.280
Rep. 3-gram -0.358 -0.248 -0.047 -0.263 -0.168 -0.519 -0.370 -0.128 -0.451 -0.299
[…]
Compression 0.358 0.069 0.026 0.516 -0.484 0.492 0.080 -0.002 0.728 -0.665
Density 0.379 0.121 0.195 0.432 -0.295 0.603 0.125 0.300 0.597 -0.474
[…]
Prom 2 8x7b -0.070 0.272 0.179 0.101 0.600 -0.112 0.366 0.248 0.077 0.767
Selene Mini 0.187 0.223 0.049 -0.112 0.684 0.231 0.291 0.088 -0.148 0.812
Qwen2.5 7B 0.080 0.122 0.086 -0.257 0.488 0.149 0.170 0.173 -0.404 0.654
Qwen2.5 72B 0.112 0.032 0.106 -0.229 0.526 0.196 0.064 0.183 -0.375 0.676
GPT-4o mini 0.660 0.387 0.319 -0.081 0.744 0.790 0.509 0.467 -0.097 0.874
GPT-4o 0.786 0.430 0.590 0.381 0.702 0.909 0.574 0.761 0.512 0.859
[…]
Table 7: Rank correlations between automatic metrics and human annotations for the Spanish data. Largest three correlations for each criteria in bold.
{tblr}
[…]
ROUGE-2 0.164 0.253 -0.064 0.037 -0.079 0.245 0.435 -0.071 0.020 -0.136
ROUGE-3 0.069 0.305 0.096 -0.016 -0.111 0.096 0.478 0.102 -0.003 -0.232
ROUGE-4 -0.005 0.284 0.139 -0.016 -0.079 0.028 0.435 0.206 -0.055 -0.186
ROUGE-L 0.491 0.263 -0.257 0.364 -0.322 0.675 0.394 -0.343 0.475 -0.479
ROUGE-su* 0.385 0.095 -0.214 0.111 -0.026 0.502 0.129 -0.290 0.179 -0.027
mROUGE-we 0.544 0.084 -0.246 0.290 -0.269 0.752 0.102 -0.310 0.460 -0.426
[…]
BertScore-r 0.449 0.116 -0.310 0.142 -0.047 0.572 0.168 -0.369 0.223 -0.091
BertScore-f 0.501 0.211 -0.139 0.480 -0.396 0.699 0.307 -0.225 0.617 -0.582
[…]
CHRF -0.121 -0.200 0.064 -0.090 0.406 -0.157 -0.233 0.131 -0.105 0.543
mMETEOR 0.026 0.137 -0.096 -0.142 0.100 0.026 0.198 -0.082 -0.186 0.083
CIDEr 0.565 0.084 -0.300 0.459 -0.417 0.786 0.114 -0.429 0.654 -0.593
[…]
Novel 1-gram -0.459 -0.326 -0.043 -0.533 0.660 -0.654 -0.418 -0.024 -0.686 0.816
Novel 2-gram -0.332 -0.411 0.011 -0.427 0.617 -0.550 -0.553 0.021 -0.484 0.785
Novel 3-gram -0.354 -0.453 -0.053 -0.364 0.575 -0.482 -0.617 -0.109 -0.437 0.737
Rep. 1-gram -0.396 -0.126 0.160 -0.364 0.100 -0.561 -0.105 0.145 -0.508 0.206
Rep. 2-gram -0.311 -0.126 0.118 -0.343 0.142 -0.478 -0.140 0.111 -0.446 0.229
Rep. 3-gram -0.311 -0.147 0.096 -0.343 0.058 -0.487 -0.149 0.072 -0.457 0.162
[…]
Compression 0.322 0.316 0.011 0.332 -0.586 0.539 0.445 0.009 0.466 -0.792
Density 0.259 0.337 0.107 0.332 -0.522 0.388 0.474 0.131 0.397 -0.682
[…]
Prom 2 8x7b -0.215 -0.128 -0.033 -0.064 0.639 -0.354 -0.194 -0.048 -0.096 0.826
Selene Mini 0.107 0.005 -0.153 0.011 0.404 0.103 -0.016 -0.202 0.038 0.561
Qwen2.5 7B 0.138 -0.106 -0.103 -0.166 0.538 0.178 -0.147 -0.088 -0.207 0.718
Qwen2.5 72B 0.160 -0.037 -0.097 -0.096 0.272 0.220 -0.050 -0.102 -0.111 0.412
GPT-4o mini 0.434 -0.116 -0.199 -0.130 0.624 0.576 -0.112 -0.280 -0.177 0.783
GPT-4o 0.587 0.032 -0.029 0.154 0.698 0.722 0.105 -0.055 0.218 0.876
[…]
Acknowledgements.
This work has been partially supported by the Basque Government (IKER-GAITU project), the Spanish Ministry for Digital Transformation and of Civil Service, and the EU-funded NextGenerationEU Recovery, Transformation and Resilience Plan (ILENIA project, 2022/TL-22/00215335 and 2022/TL22/00215334). Additional support was provided through DeepR3 (TED2021-130295B-C31) funded by MCIN/AEI/10.13039/501100011033 and European Union NextGeneration EU/PRTR; also through NL4DISMIS: Natural Language Technologies for dealing with dis- and misinformation (CIPROM/2021/021) and the grant CIBEST/2023/8, both funded by the Generalitat Valenciana.
[…]